If 86.7 g of Fe react in an excess of water, how many grams of Fe3O4 are produced?

 

3 Fe   +   4 H2O  -------->   Fe3O4     +     4 H2

1
Expert's answer
2021-05-28T00:56:20-0400

MM(Fe) = 55.84 g/mol

"n(Fe) = \frac{86.7}{55.84}=1.552 \;mol"

According to the reaction:

n(Fe3O4) "= \frac{1}{3}n(Fe) = 0.517 \;mol"

MM(Fe3O4)=231.55 g/mol

m(Fe3O4) "= 0.517 \times 231.55 = 359.5 \;g"

Answer: 359.5 g